Tohban Report 2016-12-21

From RHESSI Wiki

Jump to: navigation, search


Tohban Reports
Start Date: 14 December 2016
End Date: 21 December 2016
Tohban: Pascal Saint-Hilaire
Tohban email: Pascal@ssl.berkeley.edu
Next Tohban: Martin Fivian
List all reports



Contents

Solar Activity

Solar activity during the week was very low. Low probability of flares occurring (2% M, 6% C, according to SolarMonitor).

How many GOES flares occurred?

 Flares above B, C, M, X class were      5     0     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      1     0     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      1     0     0     0

There were RHESSI flares/GOES flares 2 / 5 over the time range 13-Dec-16 20-Dec-16


Memory Management

Despite not taking data at night (F & R), and Quiet/Vigorous decimation scheme, it was difficult to bring the memory below 50%. It is currently at 51% with thress passes remaining for the day. I.e. it is above the 50% level (start of fronts decimation) for quite a bit of time during the day.


Spacecraft Status

Detectors 3 and 9 are active and taking data. The cryocooler power was brought down to 80 W at the beginning of the week.

Data Gaps

GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2016-12-15T06:55:00.000 -- 2016-12-15T11:40:00.000 17100.000 (No solar ALMA obs during that time, from Hugh's Dec 15th email)

GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2016-12-16T04:20:00.000 -- 2016-12-16T09:00:00.000 16800.000 (No solar ALMA obs during that time, from Hugh's Dec 15th email)

...and important gaps reported on the 19th and 20th, because they have not been downloaded yet!


Detector issues

Lots of fast threshold adjustments to improve livetimes this week:


Today we increased the attenuator thresholds at account for the increase in background counts. This commanding was done today, Dec 14, 2016 at 19:36 UTC. The new attenuator thresholds are as follows:

 Command                % deadtime  Attenuator state
/IDPULOAD VALUE=0       ;   0.00      Min A0
/IDPULOAD VALUE=31      ;  12.16      Max A0
/IDPULOAD VALUE=22      ;   8.63      Min A1
/IDPULOAD VALUE=36      ;  14.12      Max A1
/IDPULOAD VALUE=20      ;   7.84      Min A2
/IDPULOAD VALUE=56      ;  21.96      Max A2
/IDPULOAD VALUE=24      ;   9.41      Min A3
/IDPULOAD VALUE=255     ; 100.00      Max A3



AGO-81837

16-351-18:41:25 start idib_chg_thrshld(3, REAR, FAST, 0xE0)
16-351-18:41:29 /IDPUDUMPTABL TABLE=DIBTBL3
16-351-18:41:29 /IDPUTABLE3 OFFSET=REARFASTDAC
16-351-18:41:50 /IDPULOAD VALUE=0xE0
16-351-18:43:39 start show_temps_avg
16-351-18:43:56 CP1: 151.3 K  
16-351-18:43:56 CP2: 150.0 K  
16-351-18:43:56 CT1: 127.5 K  
16-351-18:43:56 CT2: 125.5 K  
16-351-18:43:56 TST: 190.0 K     
16-351-18:43:56 CP1: 151.3 K  
16-351-18:43:56 CP2: 149.8 K  
16-351-18:43:56 CT1: 127.3 K  
16-351-18:43:56 CT2: 125.5 K  
16-351-18:43:56 TST: 190.2 K  
16-351-18:43:56 cp1 - ct1 = 24.33 K
16-351-18:43:56 Cryo Power: 80.0 W
16-351-18:43:56 Cryo Accel: 12.5 mG

AGO-81838

16-351-20:20:39 start idib_chg_thrshld(3,REAR,FAST,0x70)
16-351-20:20:43 /IDPUDUMPTABL TABLE=DIBTBL3
16-351-20:20:43 /IDPUTABLE3 OFFSET=REARFASTDAC
16-351-20:20:55 /IDPULOAD VALUE=0x70
16-351-20:25:51 start show_temps_avg
16-351-20:26:09 CP1: 151.2 K  
16-351-20:26:09 CP2: 150.1 K  
16-351-20:26:09 CT1: 127.7 K  
16-351-20:26:09 CT2: 126.1 K  
16-351-20:26:09 TST: 190.2 K     
16-351-20:26:09 CP1: 151.4 K  
16-351-20:26:09 CP2: 150.0 K  
16-351-20:26:09 CT1: 127.8 K  
16-351-20:26:09 CT2: 126.0 K  
16-351-20:26:09 TST: 190.3 K  
16-351-20:26:09 cp1 - ct1 = 23.36 K
16-351-20:26:09 Cryo Power: 80.7 W
16-351-20:26:09 Cryo Accel: 13.6 mG



2016-351-23:39:02 /IDPUTABLE3 OFFSET=REARFASTDAC
2016-351-23:39:13 /IDPULOAD VALUE=0xE0 ; was 0x70
2016-351-23:39:26 /IDPUTABLE3 OFFSET=FRONTFASTDAC
2016-351-23:39:32 /IDPULOAD VALUE=0x80 ; was ox60



The folloing changes have been made today at 2016-352-23:12 UTC.

Attenuator thresholds:

/IDPULOAD VALUE=0       ;   0.00  Min A0
/IDPULOAD VALUE=42      ;  16.47  Max A0
/IDPULOAD VALUE=33      ;  12.94  Min A1
/IDPULOAD VALUE=47      ;  18.43  Max A1
/IDPULOAD VALUE=31      ;  12.16  Min A2
/IDPULOAD VALUE=67      ;  26.27  Max A2
/IDPULOAD VALUE=35      ;  13.72  Min A3
/IDPULOAD VALUE=255     ; 100.00  Max A3

G3:

16-352-23:13:43 /IDPUTABLE3 OFFSET=FRONTFASTDAC
16-352-23:13:54 /IDPULOAD VALUE=0xA0 ; was 0x80

G9:

16-352-23:15:33 /IDPUTABLE9 OFFSET=FRONTFASTDAC
16-352-23:15:43 /IDPULOAD VALUE=0x80 ; was 0x50


The RHESSI read pointer was skipped forward this morning to reduce the SSR fill to ~40% in preperation for the ALMA observation. We skipped over data recorded between 351-04:22 and 351-15:12 UTC. The Front Fast thresholds were increased on both G3 and G9.

Jeremy

2016-353-17:48:43 /IDPUTABLE3 OFFSET=FRONTFASTDAC
2016-353-17:48:48 /IDPULOAD VALUE=0xB0 ; was 0xA0
2016-353-17:49:07 /IDPUTABLE9 OFFSET=FRONTFASTDAC
2016-353-17:49:16 /IDPULOAD VALUE=0x90 ; was 0x80


We had to reply the skipped data, as some of it was during an ALMA observation: (it did bring the SSR fill well below the 50% threshold, so that's good). "I was very busy today, away from email. I can tell you that those data are nearly all down, we replayed data from 2016-351-09:00 UT up to 15:13 UTC, where we skipped the read pointer to on Sunday."



2016-355-21:47:45 /IDPUTABLE3 OFFSET=REARFASTDAC
2016-355-21:47:55 /IDPULOAD VALUE=0xFF ; was 0xE0


Other notes

ALMA observations this week! Schedule, from Hugh's Dec. 15th email:

Friday AM, UT 12:30-15:30 Dec 16
2016.1.00202.S network region, 60" FOV, bands 3 and 6
Coordinates: (0,+245") at 12 UT Dec 15 with tracking for rotation -> (+220",+247") at 12 UT Dec 16
Friday PM, UT 17:00-20:00 Dec 16
2016.1.00423.S quiet Sun at disk center, 60" FOV, bands 3 and 6
Coordinates: (0,0) at 17 UT Dec 16 with tracking for rotation
Saturday AM, UT 12:30-15:30 Dec 17
2016.1.00182.S west limb at the equator for spicules, 60" FOV, bands 3+6 Coordinates: (+990",0), no rotation tracking
Sunday PM, UT 18:00-20:15 Dec 18
2016.1.00156.S prominence, west limb, 60"x120" FOV, band 3 Coordinates: (+974",+282"), off limb, no rotation tracking


Upcoming ALMA observations for next week: 12:50 UT to 20:11 UT on Thu, Sat, Sun, Tue, Wed.


Other:

The FRONT_RATIO_12_25 flag in the observing summary seems to oscillate furiously between 0 and 100%, which is odd. E.g. Check out the data from 2016/12/18.
Latest VC0 and VC1 data from http://sprg.ssl.berkeley.edu/hessiops/state-of-health/recent/vc0/welcome.shtml is from March...


Spacecraft Management

Decimation Quiet/Vigorous
HLAT Decimation Decimating
Night time data (fronts) full-time
Night time data (rears) full-time
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Both stuck in - had to adjust attenuator deadtime threshold and improve D3f livetime
Detector problems? See notes above.
Personal tools